The nitrogen at standard temperature and pressure is 1.251kgm^-3. Calculate the root mean square of nitrogen molecules.
1
Expert's answer
2019-04-18T11:34:21-0400

The  root mean square of nitrogen molecules:


"v_{rms}=\sqrt{\frac{3P}{\rho}}"

"v_{rms}=\sqrt{\frac{3(101300)}{1.251}}= 492.9\frac{m}{s}"